Write the equation of the line that contains the given point and has the given slope. (–8, –8), slope = 3
Answer by ikleyn(52949)

.
The equation of the line that contains the point (a,b) and has the slope "m" is

y - b = m*(x - a).



In your case the equation of this line is

y - (-8) = 3*(x-(-8)),  or, equivalently,

y + 8 = 3*(x+8).



You can further transform it to any other desired form.